Nine uncharged capacitors with equal capacitances

Please show the steps so I can follow along. Ill comment if incorrect.
Answer #1

The equivalent capacitance of the whole system is 1/(1/C + 1/C + 1/C + 1/C + 1/C + 1/C + 1/C + 1/C + 1/C) = 1/(9/C) = C/9
Capacitance has units of Farads or Coulombs/Volt so all we need to do is divide the amount of charge that moves through the battery during the charging process by the voltage across the battery, and that will give the total capacitance of the system:
(3.95x10^-4 C)/(7.85 V) = 5.03x10^-5 C
Now we have the total capacitance of the system, which is equal to C/9, so to solve for C:
(5.03x10^-5 C) = C/9
C = 4.527x10^-4 F